OpenStudy (anonymous):

Use the Subtraction Property of Equality to comlete the following statement: If 10x + 6 = 21, then ____ ___ = 15. A)10x-6 B)1.5 C)10x D)10x + 6 i think its A?

OpenStudy (mayankdevnani):

\[\huge 10x+6=21\] subtract -6 form both sides, we get, \[\huge 10x+6-6=21-6\] \[\huge 10x=15\] so this is our answer

OpenStudy (mayankdevnani):

understood or not? @DannyJR

OpenStudy (anonymous):

i understand the subtract 6 from both sides part but that means the answer wouldbe D?

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

@skullpatrol can yo help me out?

OpenStudy (mayankdevnani):

lol... @KingAditya you are wrong. in this question , we have to find any value which is equal to 15../ see ok... i hope you got it....

OpenStudy (anonymous):

Yeah Right.. Sorry @DannyJR.. I THOUGHT WE HAVE TO FIND THE VALUE OF "X"

OpenStudy (mayankdevnani):

Its ok...friend... @KingAditya